deidrenicoleg:

Prove the sum of two rational numbers is rational where a, b, c, and d are integers and b and d cannot be zero. Steps Reasons 1. a over b plus c over d Given 2. ad over bd plus cb over bd Multiply to get a common denominator 3. Simplify Fill in the missing step in the proof. ad times cb all over bd ad plus cb all over bd ad minus cb all over bd ad plus cb all over two times bd

deidrenicoleg:

Is it ad times cb all over bd ??

Vocaloid:

step 2: \[\frac{ ad }{ bd }+\frac{ cb }{ bd }\] for step 3: combine the numerators over one common denominator \[\frac{ ad+cb }{ bd }\]